# Analytical Expression for Continuum-continuum Transition Amplitude of Hydrogen-like Atoms with Angular-momentum Dependence

### v1

@@ Abstract
Attosecond chronoscopy typically utilises interfering two-photon transitions to access the phase information. Simulating these two-photon transitions is challenging due to the continuum-continuum transition term. The hydrogenic approximation within second-order perturbation theory has been widely used due to the existence of analytical expressions of the wave functions. So far, only (partially) asymptotic results have been derived, which fail to correctly describe the low-kinetic-energy behaviour, especially for high angular-momentum states. Here, we report an analytical expression that overcome these limitations. They are based on the Appell’s $F_{1}$ function and use the confluent hypergeometric function of the second kind as the intermediate states. We show that the derived formula can be applied to arbitrarily high angular momenta and quantitatively agrees with the numerical simulations using the time-dependent Schrödinger equation. Furthermore, we give an angular-momentum-dependent asymptotic form of the outgoing wavefunction and their continuum-continuum dipole transition amplitudes.
@@ 1 Introduction
The question “How long does the photoionsation take?” has intrigued researchers for decades ever since the early years of quantum theories. Its natural timescale of attosecond (as, ${10}^{-18}$ s) has introduced a new field of science, known as the attosecond science, which is highlighted by the 2023 Nobel Prize in Physics [1]. The last 30 years have witnessed the rapid development of the toolbox for investigating the photoionisation processes with phase information embedded. This includes, but is not limited to, high-harmonic generation (HHG) [2, 3] as a table-top XUV source thanks to the availability of the high-power lasers, the schemes of performing the phase measurement such as attosecond streaking [4, 5, 6] and the reconstruction of attosecond beating by interference of two-photon transitions (RABBIT) [7, 8], and three-dimensional momentum-resolved particle detectors such as the cold target recoil ion momentum spectrometer (COLTRIMS) [9, 10, 11], and so on. As a result, it has become possible to measure the Wigner time delay defined as [12, 13, 14]:
@@
where $\mathfrak{T}$ is the transition amplitude for photoionisation, and in the following discussions in this article, unless otherwise stated, we use the atomic units, where $e=m_{e}=\hbar=1$, and the atomic unit of time $1~{}{\rm a.u.}\approx 2.42\times{10}^{-17}~{}{\rm s}=24.2~{}{\rm as}$. Although this definition is rather straightforward, and the transition amplitudes (phases and moduli) of one-photon ionisation for atoms and molecules can be routinely calculated by e.g. random-phase approximation with exchange (RPAE) [15, 16], time-dependent local-density approximation (TDLDA) [17, 18] and Schwinger variational methods such as ePolyScat [19, 20], in reality, the schemes of the phase measurements involves two-photon transitions, which differs from one-photon ionisation. An illustration of the RABBIT measurement scheme is illustrated in figure 1. The attosecond pulse train (APT) created by HHG contains typically odd harmonics of the fundamental driving field ($\omega$), which leads to the main bands (dashed lines in figure 1). If a copy of the driving field is spatially and temporally overlapped with the APT as the dressing field, an additional sideband (solid line in figure 1) lying between the two main bands appears, which corresponds to either absorbing one photon from the lower main band (the absorption pathway), or emitting one photon from the higher main band (the emission pathway). The transition amplitudes of the absorption and the emission pathways can be written as:
@@
where $\tau_{\mathrm{XUV}}=(\phi^{\mathrm{XUV}}_{+}-\phi^{\mathrm{XUV}}_{-})/(2\omega)$ corresponds to the relative phase between the harmonic frequencies in the attosecond pulse, also known as the attochirp, and $\tau_{\rm A}=(\arg\{\mathfrak{T}^{\mathrm{(2ph)}}_{\mathrm{emi}}\}-\arg\{\mathfrak{T}^{\mathrm{(2ph)}}_{\mathrm{abs}}\})/(2\omega)$ is known as the two-photon ionisation time, since it corresponds to the phase difference of the two-photon transition amplitudes between the emission and the absorption pathways. More discussions and examples can be found in the book chapter [21].
From the theoretical point of view, the ab initio computation such as numerically solving the time-dependent Schrödinger equation (TDSE) is limited to the modelled potential with one active electron, e.g. the hydrogen-like atoms, while modelling the potential for molecules can be very challenging. A common approach for handling this problem is to separate the two-photon transition into two parts, i.e. the bound-continuum transition and the continuum-continuum transition, so that the two parts can be treated individually. This essentially stems from the lowest-order perturbation theory. The two-photon transition matrix element from an bound initial state $|n\rangle$ to a final state in the continuum with momentum $\vec{k}$ can be represented as [22]:
@@
where $Y_{l^{\prime}}^{m^{\prime}}(\theta,\varphi)$ is the spherical harmonic (in the Condon–Shortley phase convention) with angular momentum and magnetic quantum number of $l^{\prime}$ and $m^{\prime}$, respectively. For the continuum-continuum part, the transition amplitude can be expressed as the product of the angular part and the radial part:
@@
where $\langle j_{1},m_{1},j_{2},m_{2}|J,M\rangle$ is the Clebsch–Gordan coefficient and is real in the Condon–Shortley phase convention. The latter integral between the radial wavefunctions, with the $(-\pi)$-term included, is denoted as $\mathcal{T}_{l\rightarrow l^{\prime}}(k,k^{\prime})$ in the following discussion, where $k=\sqrt{2E}$ and $k^{\prime}=\sqrt{2E^{\prime}}$ are the momenta of the intermediate and final states, respectively.
Figure 1: Typical energy diagram of the RABBIT scheme. The electron is excited from the bound initial state with $\epsilon_{n}<0$ with two frequency components $\Omega_{\mathrm{abs}}$ and $\Omega_{\mathrm{emi}}$ of the attosecond pulse train (purple arrows) to the corresponding intermediate states (dashed lines) with electron kinetic energies of $E_{\mathrm{abs}}$ and $E_{\mathrm{emi}}$, respectively. The presence of the dressing field (red arrows) with frequency $\omega$ induces the continuum-continuum transition from the intermediate states to the final state (sideband) with electron kinetic energy of $E^{\prime}$ and leads to the oscillation intensity of the sideband as a function of the XUV-IR time delay $\Delta t$, as sketched by the orange wave. The partial-wave indices are denoted as $(l,m)$ for all states in the continuum, and the total transition amplitude is the superposition of the contributions from all possible pathways.
The radial wavefunctions in principle depend on the shape of the radial potential, which can be complicated for molecular cases. If we assume that the continuum-continuum transition is dominated by the long-range effect and is not sensitive to the local potential, then we can use the transition amplitudes derived from the hydrogen-like atoms. Therefore, instead of expression (1), the two-photon Wigner-like time delay is defined using the finite-difference method:
@@
which can be represented either in the spherical-coordinate basis $(\theta,\varphi)$ or in the spherical-harmonic $(l,m)$ basis. If we further assume that the continuum-continuum transition amplitude in equation (7) does not depend on $(l,m)$ or $(l^{\prime},m^{\prime})$, then the two-photon time delay can be separated into two parts:
@@
where we have approximated the derivative in expression (1) by its finite difference, and the continuum-continuum time delay is defined as
@@
where $N_{k}=\sqrt{2/(\pi k)}$ and $N_{k^{\prime}}=\sqrt{2/(\pi k^{\prime})}$ are the normalisation factors for the intermediate state and the final state, respectively. We note that the normalisation factor depends on normalising according to the cross section or normalising according to the flux. Since for the scattered wave the total flux is
@@
where $j_{\rm s}$ is the scattering flux density, $f_{\rm s}$ is the scattering amplitude, and $\mathrm{d}\sigma/\mathrm{d}\Omega$ is the differential cross section. The well-known solutions to the Coulomb wave equation are conventionally normalised according to the cross section, while the WKB approximation yields the flux-based normalisation factor, which contains an additional factor $\propto k^{-1/2}$, which is clear from comparing their asymptotic oscillation moduli [24]. The TDSE results lie typically between these two models, and they converge towards the same values for high kinetic energies [23]. A number of works have explicitly [11, 25, 26, 27, 28, 29, 30] or implicitly (e.g. to assume the two-photon time-delay difference between two species with the same or similar electron kinetic energies is equal or almost equal to their one-photon time-delay difference) [31, 32] used this assumption and have shown good agreement with the experimental results on molecules, indicating that this approximation is reasonable.
Despite the great success of this elegant model in the past decade, some features of experimental observations cannot be explained by these models. For example, the phases of these models do not depend on the angular momentum $l$ of the intermediate or the final state, which causes that the angle-dependent time delay of two-photon ionisation is simply the angle-dependent time delay of one-photon ionisation plus an isotropic continuum-continuum time delay. This can be proven as follows. Using equations (7,9), we find
@@
where the term $Y_{1}^{\mu}(\theta^{\prime},\varphi^{\prime})$ depends on the polarisation of the dressing field and does not contribute to the energy-dependent phase variation, i.e. the time delay. Hence, for helium one would expect an isotropic two-photon time delay, since its one-photon ionisation time delay has no angular dependence. This contradicts the experimental observation [33], and the mechanism was demonstrated in [34] that the moduli of the transition amplitudes for $l\rightarrow l-1$ and $l\rightarrow l+1$ have preference depending on the increase or decrease of the electron kinetic energy, which is known as the Fano’s propensity rule in the continuum-continuum transition. The same mechanism was also confirmed by a more recent experimental work from the authors [35] using the circular-XUV-circular-IR RABBIT, where the Fano’s propensity rule is embodied by the angle-dependent RABBIT phase difference between the co- and counter-rotating XUV and IR.
@@
The missing piece was partially filled by Boll et al. who replaced the final-state wavefunction from the asymptotic formula by the regular solution of the Coulomb wave equation, which is expressed by the Kummer’s function, or the confluent hypergeometric function of the first kind ${{}_{1}F_{1}}$. The result can also be written as an analytical expression using Gauss’ hypergeometric function ${{}_{2}F_{1}}$ [36]. This model quantitatively reproduces Fano’s propensity rule at high electron kinetic energies and the angle-resolved time delay of helium [36, 37]. However, the agreement deteriorates at lower energies (typically below 10 eV), particularly for larger $l^{\prime}$s, and in the same energy range the isotropic formulae proposed by Dahlström et al. also show large discrepancy compared to the TDSE. On the other hand, a lot of interesting phenomena in time delays, including shape resonance [30] and two-center interference [32] have been experimentally studied in this energy region, where the level of accuracy of the theory is partially limited by the reduced capability of the continuum-continuum delay models describing the relatively slow electrons. The general method for obtaining the $N$-photon transition amplitude from the $(N-1)$-photon transition amplitude has been known for a long time [38, 39], which involves expressing the Green’s function by the Sturmian expansion and extending into the continuum. We shall show its application as a modification of expression (5). We give its analytical expression using the Appell’s $F_{1}$ function, which quantitatively agrees with the TDSE and the experimental results in the whole energy range including the low-kinetic-energy region. We also note that the isotropic model can be extended to the $l$-dependent model by expanding the WKB approximation to the next term, and the Fano’s propensity rule and the $l$-dependent continuum-continuum transition phases can be nicely reproduced.
@@
This article is structured as following. In section 2, we first define the functions for the intermediate and final states, then we derive the analytical formula for the transition amplitudes between the Tricomi’s and Kummer’s functions. In section 3, we show multiple results including the continuum-continuum transition phases and time delays, and the Fano’s propensity rule. These results are compared to TDSE and other theoretical approaches. We also compare our method to the experiment of circular-XUV-circular-IR RABBIT on helium [35]. The conclusion and outlook can be found in section 4.
@@ 2.1 Regular and irregular solutions to the Coulomb wave equation
where $R(r)=w(r)/r$ is the radial wavefunction. Equation (19) can be adapted for arbitrary $Z>0$ by transforming $\tilde{E}=Z^{2}{E}$ and $\tilde{r}={r}/Z$, where $\tilde{E}$ and $\tilde{r}$ are the asymptotic kinetic energy and the radius, respectively. For simplicity, in the discussion below we let $Z=1$, while the treatment is general for any $Z>0$. The regular solution of the outgoing radial wavefunction that satisfies the Coulomb wave equation can be written as (note that we have absorbed the factor $k^{l+1}$ into $C_{l}(k)$, compared to the definition in [41]):
@@
We note that $f_{l}(k,r)$ and $g_{l}(k,r)$ are linear combinations of $u_{l}(k,r)$ and its complex conjugate $u_{l}^{*}(k,r)$, which are two linearly independent solutions. Asymptotically, $u_{l}(k,r)$ approaches the spherical wave with a position-dependent phase (the Coulomb phase) [41]:
@@
where $R_{\tilde{n},\tilde{l}}(\tilde{r})$ and $R_{k^{\prime},l^{\prime}}(r)$ are the radial wavefunctions of the initial (bound) and final (continuum) states, respectively, and $\mathfrak{g}_{l}^{+}(r,\tilde{r};E)$ is the Green’s function of the intermediate state that couples the initial and the final states. For the Coulomb potential, the initial and final states are the regular solutions to the Coulomb equation. The radial Green’s function in the three-dimensional space is [43, 44, 45]:
@@
where $k=\sqrt{2E}$ and $r_{<}$ ($r_{>}$) is the smaller (greater) variable between $r$ and $\tilde{r}$. The Whittaker’s functions $\mathfrak{M}_{\kappa;\mu}(z)$ and $W_{\kappa;\mu}(z)$ are related to the confluent hypergeometric functions by [40], section (13.14)
@@
If we assume that the bound-continuum transition occurs mainly in the short range due to the finite spatial extension of the initial wavefunction, and the continuum-continuum transition has major contribution from long range, we can replace $r_{>}$ and $r_{<}$ in the integral (27) by $r$ and $\tilde{r}$, respectively. Thus, the two-photon transition amplitude is proportional to the product of
@@
where for bound states $|\nu^{+}\rangle=|\nu\rangle$ and for continuum states $|\nu^{+}\rangle$ is the outgoing-wave solution corresponding to the standing-wave solution $|\nu\rangle$. The former relation is also indicated by the fact that the confluent hypergeometric functions of the first kind ($\Phi$) and second kind ($U$) truncate into the same associated Laguerre polynomials for the bound states. Compared to the approach of Boll et al. [36], integral (32) uses the intermediate of the outgoing-wave solution to the Coulomb wave equation instead of its asymptotic form, which is also used in Dahlström et al.’s approach [23]. Therefore, we propose the following calculating procedure that if the final state from one-photon ionisation has partial wave with angular momentum $l$ and asymptotically approaches $\frac{1}{r}\exp\left(\mathrm{i}kr\right)$, we choose the intermediate state as
@@
where $(l,k)$ and $(l^{\prime},k^{\prime})$ are the angular and linear momenta for the intermediate state and the final state, respectively. For the RABBIT experiment, where a photon from the dressing field is either absorbed (absorption pathway) or emitted (emission pathway), the final kinetic energy is $E^{\prime}=k^{\prime 2}/2=E~{}{\pm}^{\mathrm{abs}}_{\mathrm{emi}}~{}\omega$, while $l^{\prime}=l\pm 1$ can have both signs for each pathway, unless restricted by the magnetic quantum number in circular fields. For example, for the $1s$ electron in hydrogen, if a linearly polarised XUV photon ionises it to the $Ep_{0}$ continuum state ($z$-axis defined as the polarisation direction), and the dressing field is co-linearly polarised, both absorption and emission pathways lead to the superposition of $E^{\prime}s_{0}$ and $E^{\prime}d_{0}$ states as the final state; on the other hand, if it is ionised by a circularly polarised XUV photon and has $Ep_{+1}$ as the intermediate state ($z$-axis defined as the light propagation direction), and the dressing field is also circular with co-rotating polarisation, then for the absorption pathway only $E^{\prime}d_{+2}$ is allowed, while the emission pathway leads to the superposition of the $E^{\prime}s_{0}$ and $E^{\prime}d_{0}$ states. The dipole transition amplitude between the intermediate state and the final state is therefore $\mathcal{E}_{\omega}\mathcal{T}_{l\rightarrow l^{\prime}}(k,k^{\prime})$, and $\mathcal{T}_{l\rightarrow l^{\prime}}(k,k^{\prime})$ reads:
@@
where $N_{k}=\sqrt{2/(\pi k)}$ and $N_{k^{\prime}}=\sqrt{2/(\pi k^{\prime})}$ are the flux-based normalisation factors. In the following sections, we derive an analytical expression of equation (36) and related integrals. We will also give a simpler approximation that describes the asymptotic behaviour with consideration of the centrifugal energy following the WKB approximation.
@@
The confluent hypergeometric function of the second kind can be represented as:
with the following recurrence relations [40], equation (13.3.10):
@@
Using equation (40), we have the relation:
$\mathcal{J}^{\sigma}_{\rho}(a,b,\lambda;n^{\prime},\lambda^{\prime};Q)=\frac{1}{k}\big{[}(b-a-1)\mathcal{J}^{\sigma-1}_{\rho}(a,b-1,\lambda;n^{\prime},\lambda^{\prime};Q)$
@@
$\mathcal{J}^{1}_{\rho}(a,b,\lambda;n^{\prime},\lambda^{\prime};Q)=\frac{1}{k}\big{[}(b-a-1)\mathcal{J}^{0}_{\rho}(a,b-1,\lambda;n^{\prime},\lambda^{\prime};Q)$

### v2

@@ Abstract
Attosecond chronoscopy typically utilises interfering two-photon transitions to access the phase information. Simulating these two-photon transitions is challenging due to the continuum-continuum transition term. The hydrogenic approximation within second-order perturbation theory has been widely used due to the existence of analytical expressions of the wave functions. So far, only (partially) asymptotic results have been derived, which fail to correctly describe the low-kinetic-energy behaviour, especially for high angular-momentum states. Here, we report an analytical expression that overcome these limitations. They are based on the Appell’s $F_{1}$ function and use the confluent hypergeometric function of the second kind as the intermediate states. We show that the derived formula quantitatively agrees with the numerical simulations using the time-dependent Schrödinger equation for various angular-momentum states, which improves the accuracy compared to the other analytical approaches that were previously reported. Furthermore, we give an angular-momentum-dependent asymptotic form of the outgoing wavefunction and their continuum-continuum dipole transition amplitudes.
@@ 1 Introduction
The question “How long does the photoionisation take place?” has intrigued researchers for decades ever since the early years of quantum theories. Its natural timescale of attosecond (as, ${10}^{-18}$ s) has introduced a new field of science, known as the attosecond science, which is highlighted by the 2023 Nobel Prize in Physics [1]. The last 30 years have witnessed the rapid development of the toolbox for investigating the photoionisation processes with phase information embedded. This includes, but is not limited to, high-harmonic generation (HHG) [2, 3] as a table-top XUV source thanks to the availability of the high-power lasers, the schemes of performing the phase measurement such as attosecond streaking [4, 5, 6] and the reconstruction of attosecond beating by interference of two-photon transitions (RABBIT) [7, 8], and three-dimensional momentum-resolved particle detectors such as the cold target recoil ion momentum spectrometer (COLTRIMS) [9, 10, 11], and so on. As a result, it has become possible to measure the Wigner time delay defined as [12, 13, 14]:
@@
where $\mathfrak{T}$ is the transition amplitude for photoionisation, and in the following discussions in this article, unless otherwise stated, we use the atomic units, where $e=m_{e}=\hbar=1$, and the atomic unit of time $1~{}{\rm a.u.}\approx 2.42\times{10}^{-17}~{}{\rm s}=24.2~{}{\rm as}$. Although this definition is rather straightforward, and the transition amplitudes (phases and moduli) of one-photon ionisation for atoms and molecules can be routinely calculated by e.g. random-phase approximation with exchange (RPAE) [15, 16], time-dependent local-density approximation (TDLDA) [17, 18] and Schwinger variational methods such as ePolyScat [19, 20], in reality, the schemes of the phase measurements involve two-photon transitions, which differ from one-photon ionisations. An illustration of the RABBIT measurement scheme is illustrated in figure 1. The attosecond pulse train (APT) created by HHG contains typically odd harmonics of the fundamental driving field ($\omega$), which leads to the main bands (dashed lines in figure 1). If a copy of the driving field is spatially and temporally overlapped with the APT as the dressing field, an additional sideband (solid line in figure 1) lying between the two main bands appears, which corresponds to either absorbing one photon from the lower main band (the absorption pathway), or emitting one photon from the higher main band (the emission pathway). The transition amplitudes of the absorption and the emission pathways can be written as:
@@
where $\tau_{\mathrm{XUV}}=(\phi^{\mathrm{XUV}}_{+}-\phi^{\mathrm{XUV}}_{-})/(2\omega)$ corresponds to the relative phase between the harmonic frequencies in the attosecond pulse train, also known as the attochirp, and $\tau_{\rm A}=(\arg\{\mathfrak{T}^{\mathrm{(2ph)}}_{\mathrm{emi}}\}-\arg\{\mathfrak{T}^{\mathrm{(2ph)}}_{\mathrm{abs}}\})/(2\omega)$ is known as the two-photon ionisation time delay, since it corresponds to the phase difference of the two-photon transition amplitudes between the emission and the absorption pathways. More discussions and examples can be found in the book chapter [21].
From the theoretical point of view, the ab initio computation such as numerically solving the time-dependent Schrödinger equation (TDSE) is limited to the modelled potential with one active electron, e.g. the hydrogen-like atoms, while modelling the potential for molecules can be very challenging. A common approach for handling this problem is to separate the two-photon transition into two parts, i.e. the bound-continuum transition and the continuum-continuum (CC) transition, so that the two parts can be treated individually. This essentially stems from the lowest-order perturbation theory. The two-photon transition matrix element from an bound initial state $|n\rangle$ to a final state in the continuum with momentum $\vec{k}$ can be represented as [22]:
@@
where $Y_{l^{\prime}}^{m^{\prime}}(\theta,\varphi)$ is the spherical harmonic (in the Condon–Shortley phase convention) with angular momentum and magnetic quantum numbers of $l^{\prime}$ and $m^{\prime}$, respectively. For the continuum-continuum part, the transition amplitude can be expressed as the product of the angular part and the radial part:
@@
where $\langle j_{1},m_{1},j_{2},m_{2}|J,M\rangle$ is the Clebsch–Gordan coefficient and is real under the Condon–Shortley phase convention. The latter integral between the radial wavefunctions, with the $(-\pi)$-term included, is denoted as $\mathcal{T}_{l\rightarrow l^{\prime}}(k,k^{\prime})$ in the following discussion, where $k=\sqrt{2E}$ and $k^{\prime}=\sqrt{2E^{\prime}}$ are the momenta of the intermediate and final states, respectively.
Figure 1: Typical energy diagram of the RABBIT scheme. The electron is excited from the bound initial state with $\epsilon_{n}<0$ with two frequency components $\Omega_{\mathrm{abs}}$ and $\Omega_{\mathrm{emi}}$ of the attosecond pulse train (purple arrows) to the corresponding intermediate states (dashed lines) with electron kinetic energies of $E_{\mathrm{abs}}$ and $E_{\mathrm{emi}}$, respectively. The presence of the dressing field (red arrows) with frequency $\omega$ induces the CC transition from the intermediate states to the final state (sideband) with electron kinetic energy of $E^{\prime}$ and leads to the oscillation of the intensity of the sideband as a function of the XUV-IR time delay $\Delta t$, as sketched by the orange wave. The partial-wave indices are denoted as $(l,m)$ for all states in the continuum, and the total transition amplitude is the superposition of the contributions from all possible pathways.
The radial wavefunctions in principle depend on the shape of the radial potential, which can be complicated for molecular cases. If we assume that the CC transition is dominated by the long-range effect and is not sensitive to the local potential, then we can use the transition amplitudes derived from the hydrogen-like atoms. Therefore, instead of expression (1), the two-photon Wigner-like time delay is defined using the finite-difference method:
@@
which can be represented either in the spherical-coordinate basis $(\theta,\varphi)$ or in the spherical-harmonic $(l,m)$ basis. If we further assume that the CC transition amplitude in equation (7) does not depend on $(l,m)$ or $(l^{\prime},m^{\prime})$, then the two-photon time delay can be separated into two parts:
@@
where we have approximated the derivative in expression (1) by its finite difference, and the CC time delay is defined as
@@
where $N_{k}=\sqrt{2/(\pi k)}$ and $N_{k^{\prime}}=\sqrt{2/(\pi k^{\prime})}$ are the normalisation factors for the intermediate state and the final state, respectively. Note that the normalisation factor depends on normalising according to the cross section or normalising according to the flux. For the scattered wave the total flux is
@@
where $j_{\rm s}$ is the scattering flux density, $f_{\rm s}$ is the scattering amplitude, and $\mathrm{d}\sigma/\mathrm{d}\Omega$ is the differential cross section. The well-known solutions to the Coulomb wave equation are conventionally normalised according to the cross section, while the WKB approximation yields the flux-based normalisation factor, which contains an additional factor $\propto k^{-1/2}$, which is clear from comparing their asymptotic oscillation moduli [24]. Regarding the phase and the corresponding time delay of the CC transitions, the TDSE results lie typically between these two models, and they converge towards the same values for high kinetic energies [23]. A number of works have explicitly [11, 25, 26, 27, 28, 29, 30] or implicitly (e.g. to assume the two-photon time-delay difference between two species with the same or similar electron kinetic energies is equal or almost equal to their one-photon time-delay difference) [31, 32] used this assumption and have shown good agreement with the experimental results on molecules, indicating that this approximation is reasonable.
Despite the great success of this elegant model in the past decade, some features of experimental observations cannot be explained by these models. For example, the phases of these models do not depend on the angular momentum $l$ of the intermediate or the final state, which causes that the angle-dependent time delay of the two-photon ionisation is simply that of the one-photon ionisation plus an isotropic CC time delay. This can be proven as follows. Using equations (7) and (9), we find that
@@
where the term $Y_{1}^{\mu}(\theta^{\prime},\varphi^{\prime})$ depends on the polarisation of the dressing field and does not contribute to the energy-dependent phase variation, i.e. the time delay. Hence, for helium one would expect an isotropic two-photon time delay, since its one-photon ionisation time delay has no angular dependence. This contradicts the experimental observation [33, 34, 35], and the mechanism was demonstrated in [36] that the moduli of the transition amplitudes for $l\rightarrow l-1$ and $l\rightarrow l+1$ have preference depending on the increase or decrease of the electron kinetic energy, respectively, which is known as the Fano’s propensity rule in the CC transition. The same mechanism was also confirmed by a more recent experimental work from the authors [37] using the circular-XUV-circular-IR RABBIT, where the Fano’s propensity rule is embodied by the angle-dependent RABBIT phase difference between the co- and counter-rotating XUV and IR.
@@
The missing piece was partially filled by Boll et al. who replaced the final-state wavefunction from the asymptotic formula by the regular solution of the Coulomb wave equation, which is expressed by the Kummer’s function, or the confluent hypergeometric function of the first kind ${{}_{1}F_{1}}$. The result can also be written as an analytical expression using Gauss’ hypergeometric function ${{}_{2}F_{1}}$ [38]. This model quantitatively reproduces Fano’s propensity rule at high electron kinetic energies and the angle-resolved time delay of helium [38, 39]. However, the agreement deteriorates at lower energies (typically below 10 eV), particularly for larger $l^{\prime}$s, and in the same energy range the isotropic formulae proposed by Dahlström et al. also show large discrepancy compared to the TDSE. On the other hand, a lot of interesting phenomena in time delays, including shape resonance [30] and two-center interference [32] have been experimentally studied in this energy region, where the level of accuracy of the theory is partially limited by the reduced capability of the CC time delay models describing the relatively slow electrons. The general method for obtaining the $N$-photon transition amplitude from the $(N-1)$-photon transition amplitude has been known for a long time [40, 41], which involves expressing the Green’s function by the Sturmian expansion and extending into the continuum. We shall show its application as a modification of expression (5). We give its analytical expression using the Appell’s $F_{1}$ function, which quantitatively agrees with the TDSE and the experimental results in the whole energy range including the low-kinetic-energy region. We also note that the isotropic model can be extended to the $l$-dependent model by expanding the WKB approximation to the next term, and the Fano’s propensity rule and the $l$-dependent CC transition phases can be well reproduced.
@@
This article is structured as following. In section 2, we first define the functions for the intermediate and final states, then we derive the analytical formula for the transition amplitudes between the Tricomi’s and Kummer’s functions. In section 3, we show multiple results including the CC transition phases and time delays, and the Fano’s propensity rule. These results are compared to TDSE and other theoretical approaches. We also compare our method to the experiments of linear-linear [35] and circular-circular [37] RABBIT on helium. We briefly discuss the behaviour of high $l$-states predicted by the analytical formula. The conclusion and outlook can be found in section 4.
@@ 2.1 Regular and irregular solutions to the Coulomb wave equation
where $R(r)=w(r)/r$ is the radial wavefunction. Equation (19) can be adapted for arbitrary $Z>0$ by transforming $\tilde{E}=Z^{2}{E}$ and $\tilde{r}={r}/Z$, where $\tilde{E}$ and $\tilde{r}$ are the asymptotic kinetic energy and the radius, respectively. For simplicity, in the discussion below we let $Z=1$, while the treatment is general for any $Z>0$, and the corresponding result can be found in B. The regular solution of the outgoing radial wavefunction that satisfies the Coulomb wave equation can be written as (note that we have absorbed the factor $k^{l+1}$ into $C_{l}(k)$, compared to the definition in [43]):
@@
Note that $f_{l}(k,r)$ and $g_{l}(k,r)$ are linear combinations of $u_{l}(k,r)$ and its complex conjugate $u_{l}^{*}(k,r)$, which are two linearly independent solutions. Asymptotically, $u_{l}(k,r)$ approaches the spherical wave with a position-dependent phase (the Coulomb phase) [43]:
@@
where $R_{\tilde{n},\tilde{l}}(\tilde{r})$ and $R_{k^{\prime},l^{\prime}}(r)$ are the radial wavefunctions of the initial (bound) and final (continuum) states, respectively, and $\mathfrak{g}_{l}^{+}(r,\tilde{r};E)$ is the Green’s function of the intermediate state that couples the initial and the final states. For the Coulomb potential, the initial and final states correspond to the regular solutions to the Coulomb equation. The radial Green’s function in the three-dimensional space is [45, 46, 47]:
@@
where $k=\sqrt{2E}$ and $r_{<}$ ($r_{>}$) is the smaller (greater) variable between $r$ and $\tilde{r}$. The Whittaker’s functions $\mathfrak{M}_{\kappa;\mu}(z)$ and $W_{\kappa;\mu}(z)$ are related to the confluent hypergeometric functions according to [42], section (13.14)
@@
If we assume that the bound-continuum transition occurs mainly in the short range due to the finite spatial extension of the initial wavefunction, and the CC transition has major contribution from long range, we can replace $r_{>}$ and $r_{<}$ in the integral (27) by $r$ and $\tilde{r}$, respectively. Thus, the two-photon transition amplitude is proportional to the product of
@@
where for bound states $|\nu^{+}\rangle=|\nu\rangle$ and for continuum states $|\nu^{+}\rangle$ is the outgoing-wave solution corresponding to the standing-wave solution $|\nu\rangle$. The former relation is also indicated by the fact that the confluent hypergeometric functions of the first kind ($\Phi$) and second kind ($U$) truncate into the same associated Laguerre polynomials for the bound states. Compared to the approaches of Boll et al. [38] and Dahlström et al. [23], integral (32) uses the outgoing-wave solution to the Coulomb wave equation as the intermediate state instead of its asymptotic form. Therefore, if the final state from one-photon ionisation has partial wave with angular momentum $l$ and asymptotically approaches $\frac{1}{r}\exp\left(\mathrm{i}kr\right)$, we choose the intermediate state as
@@
where $(l,k)$ and $(l^{\prime},k^{\prime})$ correspond to the intermediate state and the final state, respectively. For the RABBIT experiment, where a photon from the dressing field is either absorbed (absorption pathway) or emitted (emission pathway), the final kinetic energy is $E^{\prime}=k^{\prime 2}/2=E~{}{\pm}^{\mathrm{abs}}_{\mathrm{emi}}~{}\omega$, while $l^{\prime}=l\pm 1$ can have both signs for each pathway, unless restricted by the magnetic quantum number in circular fields. For example, for the $1s$ electron in hydrogen, if a linearly polarised XUV photon ionises it to the $Ep_{0}$ continuum state ($z$-axis defined as the polarisation direction), and the dressing field is co-linearly polarised, both absorption and emission pathways lead to the superposition of $E^{\prime}s_{0}$ and $E^{\prime}d_{0}$ states as the final state; on the other hand, if it is ionised by a circularly polarised XUV photon and has $Ep_{+1}$ as the intermediate state ($z$-axis defined as the light propagation direction), and the dressing field is also circular with co-rotating polarisation, then for the absorption pathway only $E^{\prime}d_{+2}$ is allowed, while the emission pathway leads to the superposition of the $E^{\prime}s_{0}$ and $E^{\prime}d_{0}$ states. The dipole transition amplitude between the intermediate state and the final state is $\mathcal{E}_{\omega}\mathcal{T}_{l\rightarrow l^{\prime}}(k,k^{\prime})$, with
@@
where $N_{k}=\sqrt{2/(\pi k)}$ and $N_{k^{\prime}}=\sqrt{2/(\pi k^{\prime})}$ are the flux-based normalisation factors. In another recent work [48], Benda et al. employ a similar Green’s function method and reach an expression that is essentially the same as (36), in the context of multiphoton ionisation including the RABBIT scheme with higher-order interference pathways. The challenge for the community so far is to convert this integral into a closed-form expression, or to numerically evaluate it as accurately as possible. In the following sections, we show that equation (36) and related integrals can indeed be expressed as analytical expressions. We will also give a simpler approximation that describes the asymptotic behaviour with consideration of the centrifugal energy following the WKB approximation.
@@
The confluent hypergeometric function of the second kind has the following recurrence relations [42], equation (13.3.10):
@@
Hence, we have the relation:
$\mathcal{J}^{\sigma}_{\rho}(a,b,\lambda;n^{\prime},\lambda^{\prime};Q)=\frac{1}{\lambda}\big{[}(b-a-1)\mathcal{J}^{\sigma-1}_{\rho}(a,b-1,\lambda;n^{\prime},\lambda^{\prime};Q)$
@@
$\mathcal{J}^{1}_{\rho}(a,b,\lambda;n^{\prime},\lambda^{\prime};Q)=\frac{1}{\lambda}\big{[}(b-a-1)\mathcal{J}^{0}_{\rho}(a,b-1,\lambda;n^{\prime},\lambda^{\prime};Q)$